OpenStudy (alyssajobug):

How do I graph x-3y<6 and 2x+y>1.5

OpenStudy (opcode):

Make: < and > To equal signs. Put the equations into \(y=mx+b\) form. Then if the original equation has: > (Shade up and dotted line). < (Shade down and dotted line). Remember if you divide by a negative the sign flips!

OpenStudy (alyssajobug):

How do I find y=mx+b?

OpenStudy (anonymous):

isolate y to get y=mx+b example of a T chart to plot some points for the first one.